# ChamaleonLLM: Batch-Aware Dynamic Low-Rank Adaptation via Inference-Time Clusters

### v2

@@ Abstract
Recent advances in large language models (LLMs) have shown remarkable performance across diverse tasks. However, these models are typically deployed with fixed weights, which limits their ability to adapt dynamically to the variability inherent in real-world data during inference. This paper introduces ChamaleonLLM, a novel framework that enables inference-time adaptation of LLMs by leveraging batch-aware clustering and on-the-fly generation of low-rank updates. Unlike traditional fine-tuning approaches such as Low-Rank Adaptation (LoRA) or methods that rely on a fixed set of pre-learned uniforms (changeable masks), our method dynamically generates adaptive modifications to the decoder weights based on the aggregated statistics of clustered batches. By intelligently grouping similar inputs and computing context-aware low-rank updates via a hyper-network, ChamaleonLLM achieves significant performance gains, outperforming conventional LoRA methods while eliminating the overhead of maintaining multiple expert models. Our experiments highlight the potential of our approach to serve as a versatile and highly adaptive solution for language model inference. ChamaleonLLM is open-sourced to ensure the reproducibility of our experiments.
@@ 1 Introduction
ChamaleonLLM is designed to address this limitation by enabling the model to adapt its decoder weights during inference based on the structure and statistics of the input batch. Inputs are grouped into clusters based on their semantic and syntactic similarities. By leveraging precomputed token embeddings, the inference engine identifies coherent groups within each batch, ensuring that similar examples are processed together. Rather than relying on pre-learned and fixed uniforms (changeable masks) or static LoRA updates, ChamaleonLLM employs a hyper-network (Ha et al., 2016) to generate low-rank adaptation parameters in real-time. This dynamic generation is based on the aggregated statistics of the clustered batch, allowing the model to tailor its adaptations to the prevailing context. This approach not only removes the necessity for multiple expert models or a vast wardrobe of pre-stored masks but also leverages the collective context of the batch, leading to enhanced performance across a range of tasks. The primary contributions of this work are three-fold: (1) We introduce a method to dynamically generate low-rank updates based on batch statistics, resulting in a self-adaptive model during inference. (2) We employ a clustering-based on normalized token embeddings that groups similar inputs, ensuring context-aware adaptation. (3) We provide experimental results demonstrating the superiority of ChamaleonLLM over traditional LoRA fine-tuning.
@@ 2 Related Work
Recent research has also explored using task description embeddings derived from powerful models, such as GPT-4, to generate low-rank updates on a per-sample basis (Charakorn et al., 2024). This method generates the low-rank parameters for each input based on a task representation. While effective in certain scenarios, this per-sample adaptation may not fully exploit the shared context available when processing batches of similar inputs. ChamaleonLLM differentiates itself by using batch-level context for adaptation. By clustering inputs and computing aggregated statistics, our method generates low-rank updates that reflect the collective characteristics of the batch. Aggregating across a batch helps mitigate the noise or outlier effects in single-sample adaptation. A uniform adaptation across a cluster of similar inputs can lead to more coherent and consistent outputs. By not being limited to task-specific embeddings, our method can adapt to a broader range of scenarios, including open-domain and instruction-based tasks. Methods that use task embeddings typically generate per-sample low-rank updates based on a fixed task identifier. In contrast:
@@
To sum up, in contrast to per-sample or depth-focused adaptation methods, ChamaleonLLM exploits batch-level context to generate adaptive low-rank updates. Rather than adapting each sample independently, we cluster similar inputs and compute aggregated token embedding statistics across the batch; these statistics then drive our hyper-network to generate low-rank updates that reflect the collective context. ChamaleonLLM achieves a coherent and context-aware adaptation that naturally overcomes the limitations of fixed, per-sample updates and the storage overhead associated with large sets of pre-learned masks, such as the impact of noisy or outlier samples, and the storage overhead associated with maintaining pre-learned masks. To conclude, while traditional methods like static LoRA and changeable masks offer efficient adaptation strategies, the recent surge in dynamic approaches demonstrates a growing consensus: incorporating context—whether at the sample or task level—can significantly enhance the adaptability and robustness of LLMs. Our work builds on previous research by uniquely leveraging batch-level clustering to generate context-aware low-rank updates, thereby improving the hypernetwork-based adaptation using aggregated batch statistics.
@@ 3 Methodology
In this work, we introduce ChamaleonLLM, a dynamic adaptation framework leveraging batch-aware clustering and hyper-network–driven low-rank adaptations to efficiently tailor the model to the input context. ChamaleonLLM can be built upon any pre-trained causal language model (e.g., GPT‑2) by augmenting the conventional transformer layers and language modeling head with low-rank adaptation modules. The innovation lies in the dynamic generation of low-rank parameters based on batch-aware statistics. This process involves two major components: (1) batch-aware clustering and (2) adaptive low-rank update generation. The central hypothesis is that inputs arriving in a batch often exhibit significant similarity—whether semantic, syntactic, or stylistic. By grouping similar inputs, we can extract shared context and exploit common features in the batch. This grouping not only aids in noise reduction by averaging but also ensures that the low-rank updates are tailored to the collective characteristics of the batch rather than isolated samples. Using the normalized token embeddings, we perform k‑means clustering. The number of clusters is chosen based on the batch size to ensure each cluster is large enough to provide robust statistics while capturing meaningful variations between groups. The clustering algorithm minimizes the Euclidean distance between points and cluster centroids, and through iterative refinement, it assigns each input to a cluster that best represents its features. Once the clusters are defined, the batch is reconstructed such that each mini-batch contains inputs from the same cluster. This reorganization is critical because it ensures that subsequent low-rank adaptations are computed on a homogenous set of inputs.
@@
The core idea of dynamic adaptation is to generate low-rank modifications specific to the current batch’s context. This work proposes a more flexible alternative to LoRA that uses a hyper-network. Instead of applying a fixed set of low-rank updates (as in standard LoRA), ChamaleonLLM employs a hyper-network to produce these parameters on-the-fly. The hyper-network takes as input the mean of the token embeddings from the entire batch, processes this aggregated statistic through several fully connected layers with non-linear activations, and outputs the parameters for the low-rank update. This on-the-fly generation allows the model to adjust to the specifics of the batch dynamically, capturing nuances that a static update might miss. The overall architecture maintains the integrity of the pre-trained model while introducing minimal additional parameters. The transformer layers are wrapped with LoRA modules that operate similarly to conventional implementations, ensuring that most of the model’s capacity remains unchanged. For the LM head, the hyper-network variant ensures that adaptation is context-dependent. During inference, the model first processes the batch to compute the necessary embeddings to cluster the inputs and then generates a custom low-rank update for the LM head based on the cluster’s aggregated statistics. The final output is produced by applying this adapted LM head to the last hidden states of the transformer.
@@ 4 Experiments
For our experiments, we have used WikiText-2 (Merity et al., 2016) and Alpaca (Taori et al., 2023) datasets, providing meaningful benchmarks due to their diverse and natural language texts. The datasets are split into training and validation sets. Each text sample is tokenized using a pre-trained GPT-2 tokenizer by truncating to a maximum length and padding to ensure uniform sequence lengths. LM input token embeddings are computed for every example, and normalized to facilitate robust clustering. For the Alpaca dataset, we calculate the token embeddings from the instruction part of the input prompts. A crucial part of ChamaleonLLM is creating data loaders that reflect the clustered nature of the inputs. Using the precomputed token embeddings, we apply k‑means clustering with the number of clusters based on the desired batch size and overall dataset size. After clustering, indices are grouped so each mini-batch contains examples from a single cluster. This ensures that each batch is contextually coherent. We implement a custom data loader that leverages these clusters as batch samplers. The data loader returns batches fed into the model during training and evaluation. The evaluation is performed at the end of each epoch on the validation set, where batches are processed through the clustering pipeline, and the average loss is computed. This systematic evaluation ensures that our comparisons are fair and reflect real-world performance. All codes for ChamaleonLLM are open-sourced to ensure the reproducibility of experimental results.
@@
  • ChamaleonLLM Fine-Tuning: In this regime, while the transformer layers continue to use static LoRA modules, the LM head is adapted by a hyper-network LoRA module that generates low-rank update parameters based on the mean token embeddings of the batch. The loss is computed on the LM head outputs, and optimized similarly.
Experimental results demonstrate that ChamaleonLLM significantly improves over the traditional LoRA approach, consistently achieving a lower average validation loss on both datasets, and suggest that dynamic adaptation based on batch statistics leads to better convergence and generalization. As shown in Table 1, ChamaleonLLM achieves a validation loss reduction of approximately 25% compared to traditional LoRA, along with a perplexity improvement of roughly 12%. Table 2 further indicates that it reduces the validation loss by almost 30% on instruction fine-tuning. These improvements underscore the effectiveness of the batch-level context adaptation strategy and can largely be attributed to the exploitation of batch-level context. By averaging token embeddings over each batch, the hyper-network input becomes less sensitive to individual outliers, leading to robust low-rank parameter generation. Clustering groups together semantically or stylistically similar inputs allows the hyper-network to extract stronger, more relevant signals for adaptation.
Although ChamaleonLLM has more trainable parameters than LoRA, this increase is justified by many practical benefits, as it dynamically generates low-rank updates on-the-fly rather than training and storing separate adaptation matrices for each sample or task. Avoiding the storage of numerous pre-learned matrices significantly reduces memory and processing overheads during inference. The additional parameters enable context-aware, batch-level adaptation, leading to better performance. Despite it also introduces some extra computational steps during inference, the overhead is modest compared to the overall inference time. The trade-off between increased computation and improved performance is highly favorable. Our approach is not tied to pre-defined task categories and adapts to the nuances in the data at inference time, offering a versatile solution for open-domain tasks.
@@ 5 Conclusion
In this paper, we presented ChamaleonLLM—a novel framework that enables inference-time adaptation of large language models through batch-aware clustering and dynamic low-rank parameter generation. By clustering similar inputs and using a hyper-network to generate low-rank updates based on the aggregated batch statistics, our method provides a flexible and efficient alternative to static fine-tuning approaches such as traditional LoRA or uniform mask blending. Our extensive experiments on WikiText-2 and Alpaca demonstrate that ChamaleonLLM achieves lower validation loss and perplexity than baseline methods with dynamic batch-aware contextual adaptation. The proposed framework reduces the need for storing multiple expert masks and adapts to diverse input distributions in real-time, making it highly suitable for modern, dynamic inference environments.

### v3

# ChameleonLLM: Batch-Aware Dynamic Low-Rank Adaptation via Inference-Time Clusters

@@ Abstract
Recent advances in large language models (LLMs) have shown remarkable performance across diverse tasks. However, these models are typically deployed with fixed weights, which limits their ability to adapt dynamically to the variability inherent in real-world data during inference. This paper introduces ChameleonLLM, a novel framework that enables inference-time adaptation of LLMs by leveraging batch-aware clustering and on-the-fly generation of low-rank updates. Unlike traditional fine-tuning approaches such as Low-Rank Adaptation (LoRA) or methods that rely on a fixed set of pre-learned uniforms (changeable masks), our method dynamically generates adaptive modifications to the decoder weights based on the aggregated statistics of clustered batches. By intelligently grouping similar inputs and computing context-aware low-rank updates via a hyper-network, ChameleonLLM achieves significant performance gains, outperforming conventional LoRA methods while eliminating the overhead of maintaining multiple expert models. Our experiments highlight the potential of our approach to serve as a versatile and highly adaptive solution for language model inference. ChameleonLLM is open-sourced to ensure the reproducibility of our experiments.
@@ 1 Introduction
ChameleonLLM is designed to address this limitation by enabling the model to adapt its decoder weights during inference based on the structure and statistics of the input batch. Inputs are grouped into clusters based on their semantic and syntactic similarities. By leveraging precomputed token embeddings, the inference engine identifies coherent groups within each batch, ensuring that similar examples are processed together. Rather than relying on pre-learned and fixed uniforms (changeable masks) or static LoRA updates, ChameleonLLM employs a hyper-network (Ha et al., 2016) to generate low-rank adaptation parameters in real-time. This dynamic generation is based on the aggregated statistics of the clustered batch, allowing the model to tailor its adaptations to the prevailing context. This approach not only removes the necessity for multiple expert models or a vast wardrobe of pre-stored masks but also leverages the collective context of the batch, leading to enhanced performance across a range of tasks. The primary contributions of this work are three-fold: (1) We introduce a method to dynamically generate low-rank updates based on batch statistics, resulting in a self-adaptive model during inference. (2) We employ a clustering-based on normalized token embeddings that groups similar inputs, ensuring context-aware adaptation. (3) We provide experimental results demonstrating the superiority of ChameleonLLM over traditional LoRA fine-tuning.
@@ 2 Related Work
Recent research has also explored using task description embeddings derived from powerful models, such as GPT-4, to generate low-rank updates on a per-sample basis (Charakorn et al., 2024). This method generates the low-rank parameters for each input based on a task representation. While effective in certain scenarios, this per-sample adaptation may not fully exploit the shared context available when processing batches of similar inputs. ChameleonLLM differentiates itself by using batch-level context for adaptation. By clustering inputs and computing aggregated statistics, our method generates low-rank updates that reflect the collective characteristics of the batch. Aggregating across a batch helps mitigate the noise or outlier effects in single-sample adaptation. A uniform adaptation across a cluster of similar inputs can lead to more coherent and consistent outputs. By not being limited to task-specific embeddings, our method can adapt to a broader range of scenarios, including open-domain and instruction-based tasks. Methods that use task embeddings typically generate per-sample low-rank updates based on a fixed task identifier. In contrast:
@@
To sum up, in contrast to per-sample or depth-focused adaptation methods, ChameleonLLM exploits batch-level context to generate adaptive low-rank updates. Rather than adapting each sample independently, we cluster similar inputs and compute aggregated token embedding statistics across the batch; these statistics then drive our hyper-network to generate low-rank updates that reflect the collective context. ChameleonLLM achieves a coherent and context-aware adaptation that naturally overcomes the limitations of fixed, per-sample updates and the storage overhead associated with large sets of pre-learned masks, such as the impact of noisy or outlier samples, and the storage overhead associated with maintaining pre-learned masks. To conclude, while traditional methods like static LoRA and changeable masks offer efficient adaptation strategies, the recent surge in dynamic approaches demonstrates a growing consensus: incorporating context—whether at the sample or task level—can significantly enhance the adaptability and robustness of LLMs. Our work builds on previous research by uniquely leveraging batch-level clustering to generate context-aware low-rank updates, thereby improving the hypernetwork-based adaptation using aggregated batch statistics.
@@ 3 Methodology
In this work, we introduce ChameleonLLM, a dynamic adaptation framework leveraging batch-aware clustering and hyper-network–driven low-rank adaptations to efficiently tailor the model to the input context. ChameleonLLM can be built upon any pre-trained causal language model (e.g., GPT‑2) by augmenting the conventional transformer layers and language modeling head with low-rank adaptation modules. The innovation lies in the dynamic generation of low-rank parameters based on batch-aware statistics. This process involves two major components: (1) batch-aware clustering and (2) adaptive low-rank update generation. The central hypothesis is that inputs arriving in a batch often exhibit significant similarity—whether semantic, syntactic, or stylistic. By grouping similar inputs, we can extract shared context and exploit common features in the batch. This grouping not only aids in noise reduction by averaging but also ensures that the low-rank updates are tailored to the collective characteristics of the batch rather than isolated samples. Using the normalized token embeddings, we perform k‑means clustering. The number of clusters is chosen based on the batch size to ensure each cluster is large enough to provide robust statistics while capturing meaningful variations between groups. The clustering algorithm minimizes the Euclidean distance between points and cluster centroids, and through iterative refinement, it assigns each input to a cluster that best represents its features. Once the clusters are defined, the batch is reconstructed such that each mini-batch contains inputs from the same cluster. This reorganization is critical because it ensures that subsequent low-rank adaptations are computed on a homogenous set of inputs.
@@
The core idea of dynamic adaptation is to generate low-rank modifications specific to the current batch’s context. This work proposes a more flexible alternative to LoRA that uses a hyper-network. Instead of applying a fixed set of low-rank updates (as in standard LoRA), ChameleonLLM employs a hyper-network to produce these parameters on-the-fly. The hyper-network takes as input the mean of the token embeddings from the entire batch, processes this aggregated statistic through several fully connected layers with non-linear activations, and outputs the parameters for the low-rank update. This on-the-fly generation allows the model to adjust to the specifics of the batch dynamically, capturing nuances that a static update might miss. The overall architecture maintains the integrity of the pre-trained model while introducing minimal additional parameters. The transformer layers are wrapped with LoRA modules that operate similarly to conventional implementations, ensuring that most of the model’s capacity remains unchanged. For the LM head, the hyper-network variant ensures that adaptation is context-dependent. During inference, the model first processes the batch to compute the necessary embeddings to cluster the inputs and then generates a custom low-rank update for the LM head based on the cluster’s aggregated statistics. The final output is produced by applying this adapted LM head to the last hidden states of the transformer.
@@ 4 Experiments
For our experiments, we have used WikiText-2 (Merity et al., 2016) and Alpaca (Taori et al., 2023) datasets, providing meaningful benchmarks due to their diverse and natural language texts. The datasets are split into training and validation sets. Each text sample is tokenized using a pre-trained GPT-2 tokenizer by truncating to a maximum length and padding to ensure uniform sequence lengths. LM input token embeddings are computed for every example, and normalized to facilitate robust clustering. For the Alpaca dataset, we calculate the token embeddings from the instruction part of the input prompts. A crucial part of ChameleonLLM is creating data loaders that reflect the clustered nature of the inputs. Using the precomputed token embeddings, we apply k‑means clustering with the number of clusters based on the desired batch size and overall dataset size. After clustering, indices are grouped so each mini-batch contains examples from a single cluster. This ensures that each batch is contextually coherent. We implement a custom data loader that leverages these clusters as batch samplers. The data loader returns batches fed into the model during training and evaluation. The evaluation is performed at the end of each epoch on the validation set, where batches are processed through the clustering pipeline, and the average loss is computed. This systematic evaluation ensures that our comparisons are fair and reflect real-world performance. All codes for ChameleonLLM are open-sourced to ensure the reproducibility of experimental results.
@@
  • ChameleonLLM Fine-Tuning: In this regime, while the transformer layers continue to use static LoRA modules, the LM head is adapted by a hyper-network LoRA module that generates low-rank update parameters based on the mean token embeddings of the batch. The loss is computed on the LM head outputs, and optimized similarly.
Experimental results demonstrate that ChameleonLLM significantly improves over the traditional LoRA approach, consistently achieving a lower average validation loss on both datasets, and suggest that dynamic adaptation based on batch statistics leads to better convergence and generalization. As shown in Table 1, ChameleonLLM achieves a validation loss reduction of approximately 25% compared to traditional LoRA, along with a perplexity improvement of roughly 12%. Table 2 further indicates that it reduces the validation loss by almost 30% on instruction fine-tuning. These improvements underscore the effectiveness of the batch-level context adaptation strategy and can largely be attributed to the exploitation of batch-level context. By averaging token embeddings over each batch, the hyper-network input becomes less sensitive to individual outliers, leading to robust low-rank parameter generation. Clustering groups together semantically or stylistically similar inputs allows the hyper-network to extract stronger, more relevant signals for adaptation.
Although ChameleonLLM has more trainable parameters than LoRA, this increase is justified by many practical benefits, as it dynamically generates low-rank updates on-the-fly rather than training and storing separate adaptation matrices for each sample or task. Avoiding the storage of numerous pre-learned matrices significantly reduces memory and processing overheads during inference. The additional parameters enable context-aware, batch-level adaptation, leading to better performance. Despite it also introduces some extra computational steps during inference, the overhead is modest compared to the overall inference time. The trade-off between increased computation and improved performance is highly favorable. Our approach is not tied to pre-defined task categories and adapts to the nuances in the data at inference time, offering a versatile solution for open-domain tasks.
@@ 5 Conclusion
In this paper, we presented ChameleonLLM—a novel framework that enables inference-time adaptation of large language models through batch-aware clustering and dynamic low-rank parameter generation. By clustering similar inputs and using a hyper-network to generate low-rank updates based on the aggregated batch statistics, our method provides a flexible and efficient alternative to static fine-tuning approaches such as traditional LoRA or uniform mask blending. Our extensive experiments on WikiText-2 and Alpaca demonstrate that ChameleonLLM achieves lower validation loss and perplexity than baseline methods with dynamic batch-aware contextual adaptation. The proposed framework reduces the need for storing multiple expert masks and adapts to diverse input distributions in real-time, making it highly suitable for modern, dynamic inference environments.